OpenStudy (anonymous):

Rosalyn works no more than 20 hours a week during the school year. She is paid $10 an hour for tutoring geometry students and $7 an hour for delivering pizzas for Pizza King. She wants to spend at least 3 hours but no more than 8 hours a week tutoring. Find Rosalyn's maximum earnings.

OpenStudy (anonymous):

hey..ur name maninblack is it reffering to the Princess Bride?

OpenStudy (anonymous):

no Goodman, Roy Orbison sort of, but mainly I like wearing black :-)

OpenStudy (anonymous):

oh.srry i was reading that earlier and there is a man in black in it...anyways i need help

OpenStudy (anonymous):

is there more to this question?

OpenStudy (anonymous):

well..i dunno if u've been through the finding minimums and maximums before, but thats what it is about

OpenStudy (anonymous):

if not, if she does 8 hours tutoring = 80 plus, 12 hours pizza delivery = 84 total 164 otherwise 3 hours tutoring = 30 and 17 hours pizza = 119 total 149

OpenStudy (anonymous):

so 149 is the maximum?

OpenStudy (anonymous):

164 if she does 8 hours tutoring 149 if she does 3 hours tutoring and the rest is pizza delivery

OpenStudy (anonymous):

im kinda confused...

OpenStudy (anonymous):

ok if she wants to tutor for 8 hours and deliver pizza for the rest of her 20 hours of work = $10 * 8 hours = $80 plus $7 times 12 hours = $84 add the to totals = $164 from both jobs

OpenStudy (anonymous):

thats the maximum she can earn

OpenStudy (anonymous):

ohhhhh!!! now i get it!! thnx a lot!!

OpenStudy (anonymous):

if she only works 3 hours tutoring thats $10 * 3 hours = $30 plus $7 * 17 hours = $119 to give a total of $149 minimum she can earn

OpenStudy (anonymous):

as long as she doesn't slack off and go to the beach :-P

OpenStudy (anonymous):

LOL...thnx anyways!

OpenStudy (anonymous):

enjoy if you can

OpenStudy (anonymous):

;D